If 5 hens can lay 24 eggs in 5 days how many days are needed for 8 hens to lay 20 eggs
Answer by stanbon(75887)

If 5 hens can lay 24 eggs in 5 days how many days are needed for 8 hens to lay 20 eggs
----
# of days is directly related to number of eggs and indirectly
related to # of hens.
---
d = k*e/h
----------
Solve for "k" using "5 hens can lay 24 eggs in 5 days".
5 = k*24/5
k = 24/25
-------
Equation:
d = (24/25)e/h
------
How many days are needed for 8 hens to lay 20 eggs?
d = (24/25)*20/8
----
# of days = (60/25) = 12/5 = 2.4 days
======
Cheers,
Stan H.